OpenStudy (anonymous):

in 1991 the life expectancy of males in a certain country was 64.6 years . in 1995 it was 68.4 years. let E represent the life expectancy in year T represent the number of years since 1991 the liner function E(t) that fits data is e(t)=__ t+__ round to the nearest 10th use the function to predict the life expectancy of males in 2005 E(14)= round to the nearest 10th

OpenStudy (anonymous):

you're trying to find a line that goes through (64.6,0) and (68.4,4) use e = mt + c m = slope = 4/(68.4 - 64.6) = 4/3.8

OpenStudy (saifoo.khan):

\begin{array}l\color{#7a284a}{\text{u}}\color{#983152}{\text{ }}\color{#73a259}{\text{a}}\color{#2a75a9}{\text{l}}\color{#743378}{\text{r}}\color{#625952}{\text{e}}\color{#449a28}{\text{a}}\color{#85a499}{\text{d}}\color{#94894a}{\text{y}}\color{#6a8115}{\text{ }}\color{#254858}{\text{g}}\color{#156a48}{\text{o}}\color{#427282}{\text{t}}\color{#67424a}{\text{.}}\color{#313745}{\text{}}\end{array}

OpenStudy (saifoo.khan):

=)

OpenStudy (anonymous):

huh